OpenStudy (anonymous):

@dumbcow help ? Find the first, fourth, and tenth terms of the arithmetic sequence described by the given rule. a(n)=-3+(n-1)(-2.2)

OpenStudy (dumbcow):

well since they already give you the rule plug in n = 1,4,10 and see what you get for each one

OpenStudy (anonymous):

–3, –9.6, –22.8 –2.2, –11.8, –19.8 –3, –11.8, –25 0, –6.6, –19.8

OpenStudy (dumbcow):

lol yes those are answer choices what do you get when you plug in n=1, n=4, and n=10

OpenStudy (anonymous):

how do you plug it in i think my method is wrong

OpenStudy (dumbcow):

ok if n=1 , then n-1 = 1-1 = 0 right

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

-3 + 0*-2.2 = ?

OpenStudy (anonymous):

i had that before okay so im right do i have to divide

OpenStudy (dumbcow):

hmm no theres no division sign multiply by 0, what do you get?

OpenStudy (anonymous):

youll get zero

OpenStudy (dumbcow):

yep -3 + 0 = -3

OpenStudy (anonymous):

okay

OpenStudy (dumbcow):

now try n=4 step 1: subtract 1 step 2: multiply by -2.2 step 3: add -3

OpenStudy (anonymous):



OpenStudy (dumbcow):